2 C2H6(g)          +        2 O2(g) =  CO2(g)    +        2 H2O(I)

1.   How many grams of H2O will be produced by 58.2 L of CH4 at STP? Assume an excess of O2.


1
Expert's answer
2021-10-15T12:32:24-0400

Molar mass of CH4= 16.04 g/mol

58.2/22.4= 2.598moles

Molar mass of H2O= 18.01528 g/mol


Mole ratio 2:2

= 2.598×18.01528= 46.81g